OpenStudy (anonymous):

Mika bought 3L of milk for $6. Each bottle is 1L. How much did she pay for 1L?

OpenStudy (anonymous):

Ratio problem: \[\frac{3}{6} = \frac{1}{x}\] Solve for x: \[\frac{3x}{6}=\frac{1x}{x}\] \[\frac{3x}{6} = 1\] \[\frac{3x\times6}{6}=1\times6\] \[3x = 6\] \[\frac{3x}{3} = \frac{6}{3}\] \[x = 2\]

OpenStudy (anonymous):

thanks

OpenStudy (doc.brown):

Or conversion problem\[1L\times(\frac{$6}{3L})\]\[1\cancel{L}\times(\frac{$6}{3\cancel{L}})\]\[\frac{$6}{3}=$2\]

OpenStudy (anonymous):

This is a unit rate/unit price problem. The "Unit Price" (or "unit cost") tells you the cost per liter, per kilogram, per pound, etc, of what you want to buy. So, all you really have to do is turn 3L of milk for $6 into a fraction. When doing unit prices, I usually put the price as the numerator and the number of items as the denominator, but it doesn't really matter. So, I come up with \[\frac{ 6 }{ 3 }\] Now, I must reduce 6/3 into simplest form since 6/3 can be reduced. I must find the LCD, or the Least Common Denominator, which is 3 in this case, because 3 can go into both 3 and 6. So, 3 goes into 3 once, so the denominator is 1 and 3 goes into 6 twice. So 2 is the numerator. \[3\div3=1\] \[3\div6=2\] so we have, \[\frac{ 2 }{ 1 }\] which we know is just 2 wholes or just 2 because any number over 1 in a fraction is a whole number. c: